# ThinkVideo: High-Quality Reasoning Video Segmentation with Chain of Thoughts

### v2

@@ Abstract
Reasoning Video Object Segmentation is a challenging task, which generates a mask sequence from an input video and an implicit, complex text query. Existing works probe into the problem by finetuning Multimodal Large Language Models (MLLM) for segmentation-based output, while still falling short in difficult cases on videos given temporally-sensitive queries, primarily due to the failure to integrate temporal and spatial information. In this paper, we propose ThinkVideo, a novel framework which leverages the zero-shot Chain-of-Thought (CoT) capability of MLLM to address these challenges. Specifically, ThinkVideo utilizes the CoT prompts to extract object selectivities associated with particular keyframes, then bridging the reasoning image segmentation model and SAM2 video processor to output mask sequences. The ThinkVideo framework is training-free and compatible with closed-source MLLMs, which can be applied to Reasoning Video Instance Segmentation. We further extend the framework for online video streams, where the CoT is used to update the object of interest when a better target starts to emerge and becomes visible. We conduct extensive experiments on video object segmentation with explicit and implicit queries. The results show that ThinkVideo significantly outperforms previous works in both cases, qualitatively and quantitatively. Code will be released at github.com/DanielSHKao/ThinkVideo.
Figure 1: ThinkVideo is a novel framework for Reasoning Video Segmentation utilizing the zero-shot Chain-of-Thought (CoT) capability of pretrained multimodal Large Language Models, and correctly segments the fast-moving player before, during, and after his three-point shot after CoT, given a temporally sensitive query “Which player makes a three-point shot in this basketball game” which requires both spatial and temporal reasoning. Refer to our supplemental videos for showcase demos.
@@ 1 Introduction
Reasoning Video Object Segmentation (Reasoning VOS) presents a formidable challenge in the field of computer vision, requiring the generation of a mask sequence from an input video alongside an implicit and often complex text query [82, 4]. Unlike traditional vision-language tasks that directly associate visual data with textual descriptions, Reasoning VOS requires advanced cognitive capabilities given the implicit nature of the queries. This task is particularly demanding due to the dynamic nature of video data, where factors such as temporally sensitive queries, occlusions, or rapid object movements can complicate the segmentation process.
@@
Existing approaches of reasoning segmentation focus on finetuning the multimodal Large Language Models (MLLMs) for segmentation-based tokens. While achieving good results in Reasoning VOS, they require time-consuming finetuning process and still suffer from low accuracy when temporal reasoning is required. An implicit and time sensitive query, “Which player makes a three-point shot in this basketball game?”, for instance, is more challenging than “Which player is wearing black?” (Fig. 1). A query is time or temporally sensitive if the relevant video objects with the specified action context (e.g., three-point shot), do not appear uniformly over time but rather are concentrated at certain intervals, requiring high-level temporal reasoning to correctly find them to the satisfaction of the query requirement. Even though reasoning segmentation has achieved impressive performance on images [34, 31, 41] and videos [4], current results are far from perfect in the video domain, showcasing failure in integrating temporal information with spatial and textual context in satisfying complex and time-sensitive queries.
To address these challenges, we introduce ThinkVideo, an innovative framework designed to leverage the zero-shot Chain-of-Thought (CoT) capabilities of pre-trained MLLMs (Fig. 1). Instead of directly predicting the mask tokens, ThinkVideo breaks down the reasoning process into a sequence of “thinking” steps with multiple synthetic pairs of questions and answers. This thinking procedure will generate instance-level selectivities linked to particular keyframes, then bridging the reasoning image segmentation model with the SAM2 [60] video processor to generate coherent mask sequences. This approach not only enhances the model’s ability to reason about object relationships and movements but also allows for more accurate segmentation in complex video environments.
@@
One of the standout features of the ThinkVideo framework is its training-free nature, making it compatible with open-source and closed-source MLLMs. We leverage the state-of-the-art Gemma3 [65] and GPT-4o [29] as the “thinking” agent in our ThinkVideo framework to generate keyframe selectivity. To evaluate the effectiveness of ThinkVideo, we conduct comprehensive experiments on referring VOS datasets (e.g., MeViS [20] and Refer-DAVIS-17 [58]) and reasoning VOS datasets (e.g., ReVOS [82] and ReasonVOS [4]). We demonstrate that ThinkVideo significantly outperforms existing referring and reasoning VOS approaches, qualitatively and quantitatively, with its more general formulation based on Reasoning Video Instance Segmentation (Reasoning VIS), which can output multiple instance-level mask sequences that satisfy the user query. In addition, we propose a variant version of ThinkVideo for online video streams, where CoT is used to update the target object when a new object starts to appear in the middle of the video that aligns better with the text query.
In conclusion, our contributions include:
  • We propose ThinkVideo, a zero-shot-CoT framework for Reasoning VIS and VOS.
  • ThinkVideo significantly outperforms existing SOTA qualitatively and quantitatively.
  • We extend ThinkVideo to process online video streams.
@@ 2 Related Work
Video segmentation. Video object segmentation (VOS) tracks objects in a given video and output corresponding mask sequences. Early VOS approaches [47, 10] rely on online first-frame finetuning, thereby constrained to specific applications because of the high time cost. To speed up inference, follow-up works adopt efficient online learning algorithm [61, 7], MRF graph inference [5], temporal CNN [81, 27], capsule routing [24], tracking [30, 68, 57, 12], embedding learning [83, 66, 13] and space-time matching [28, 54, 16]. Recent VOS methods allow various control modalities, such as categories [14, 70, 78, 89, 90, 93], points [60, 59, 26], segmentation masks [15, 17, 55, 57, 60], or text query [8, 20, 75, 62, 84, 37, 76, 48, 18], making the segmentation more flexible and interactive.
@@
Specifically, building on the advanced development of LLM, Reasoning VOS focuses on generating mask sequences from a highly implicit and complex text query. VISA [82] first selects a keyframe and follows the core idea of LISA [34] by generating a segmentation-oriented token with finetuned image-based MLLM. Video-LISA [4] and HyperSeg [71] encode the video input in a multi-grained manner and concatenate the text query, feeding both to the LLM for segmentation-based outputs. However, none of the prior works incorporate adequate consideration of CoT reasoning ability for spatial and temporal information in Reasoning VOS. Our ThinkVideo is the first zero-shot-CoT framework compatible with pre-trained MLLM and image segmentation module, making it flexible to a boarder audience.
@@
Chain-of-Thought MLLM. Drawing inspiration from the exceptional reasoning capabilities of Large Language Models (LLMs), Multi-modal Large Language Models (MLLMs) seek to integrate visual components to enhance multimodal perception and reasoning. These models are often employed to produce captions or respond to inquiries based on visual and textual inputs, typically achieved by techniques such as prompt engineering [42, 63, 73], cross-attention mechanism [1], Q-Former [19, 36], prompt tuning [88], linear projection [32, 39], and unified model architectures [56, 80]. These MLLMs have greatly revolutionized the traditional vision-centric tasks like image segmentation [34, 79, 87, 25, 6], object detection [87, 86, 85, 11, 46], video segmentation [82, 4, 71], and object tracking [67, 53, 94].
@@
In addition to these advancements, CoT-based LLMs have been proposed to improve the reasoning capability [72, 69]. Existing CoT approaches tend to concentrate on inference through progressive reasoning demonstrations [91, 72, 45] or by utilizing explicit prompts that clearly outline each step [33]. However, enabling CoT capabilities for MLLMs typically undergo fine-tuning on specifically designed multimodal CoT datasets [52, 92, 43] or incorporate intricate intermediate representations [51, 64], which restricts their usability and scalability due to data accessibility and time complexity. Closer to our work, recent works, such as ThinkFirst [31] and Seg-Zero [41], have successfully incorporate CoT MLLM in reasoning image segmentation, while still falling short of “thinking” in the temporal domain. In contrast, we aim to, for the first time, abridge this gap by exploring zero-shot CoT reasoning of pre-trained MLLMs and apply them in the area of video segmentation.
3 ThinkVideo: Multi-Agent Framework
@@
ThinkVideo is a multi-agent framework that consists of three network components: an MLLM keyframe selector $\mathcal{F}_{key}$, a reasoning image segmentation model $\mathcal{F}_{seg}$, and a video processor $\mathcal{F}_{vid}$, see Fig. 2.
Figure 2: ThinkVideo for Reasoning VIS where Reasoning VOS is a special case.
@@
Input and Output Settings. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$ that can be very complex, ThinkVideo is a Reasoning Video Instance Segmentation model $\psi_{\theta}(\cdot)$ to generate a list of non-overlapping binary mask sequences $\hat{y}=\left\{\mathcal{M}_{i}\in\mathbb{R}^{T\times H\times W}\right\}_{i=1}^{k}$:
@@
where $k$ indicates the number of object instances of interest that satisfy the requirements of input query $q$. Reasoning Video Object Segmentation (VOS) is clearly reducible to Reasoning Video Instance Segmentation (VIS) by setting $k=1$.
@@
Query requirement. ThinkVideo shares a similar query setting with the first Reasoning Video Object Segmentation [82], where the input query $q$ is complex or implicit (e.g., “Please find the vehicle in the video which can accommodate at least four people.”), in contrast to the Referring VOS [62] task which takes explicit and direct query (e.g. “Please segment the red car in the video.”).
@@ 3.1 Multi-Agentic Framework
Figure 3: Illustration of the CoT Process. The MLLM agent is prompted to synthesize chain-of-thought questions and answers. This CoT process enhances the in-depth temporal and spatial understanding for the keyframe candidates. Based on the CoT result, the MLLM agent outputs an instance-level object list, which contains the object-specific keyframe and description within the frame. Refer to the supplementary materials for the detailed prompt and output examples.
MLLM Keyframe selector. The keyframe selector is an MLLM agent, e.g., GPT-4o [29] or Gemma3 [65], which generates keyframe selectivites from a list of keyframe candidates, denoted as $\{I^{c}_{i}\}$, with a CoT thinking process. The ultimate goal of this MLLM agent is to find all the object instances that meet the query requirement and select their respective keyframes, where the instance can be observed in that frame with the least effort. The output contains a list of selected object instances, each of which corresponds to an object index, a respective keyframe, and a keyframe-related object description (e.g., “the player in black jersey at the right side of the frame” ).
To achieve this goal, the agent is prompted to automatically synthesize a series of questions for each keyframe candidate and answer them one-by-one. Starting from more general information like “What can be seen in the frame?” and eventually ending with detailed questions like “Are there any other objects that meet the query requirement?”, this CoT process aims to provide in-depth spatial and temporal reasoning for the agent to find the correct objects of interest.
After analyzing all the keyframe candidates, the agent will generate an output list composed of all objects of interest and the corresponding keyframes based on its progressive question-and-answer pairs. This process can be visualized in Fig. 3.
@@
3.2 ThinkVideo: Reasoning Video Instance Segmentation
CoT can be applied to the entire video when available for reasoning instance segmentation. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$, we first uniformly sample a number of keyframe candidates $\{I^{c}_{i}\in x_{v}\}_{i=1}^{T^{\prime}}$, where $T^{\prime}=\lfloor T/\xi\rfloor$ for some $\xi<T$. Then we send these candidates and the user query $q$ to the keyframe selector $\mathcal{F}_{key}$ to generate the chain-of-thought keyframe selectivity $S$, which is a list of instance-level text descriptions $\{s_{i}\}_{i=1}^{k}$ with respect to their corresponding keyframes $\{f_{i}\}_{i=1}^{k}$ (e.g., “the red car” in $I^{c}_{2}$, “the blue motorcycle” in $I^{c}_{4}$, …), where $k$ is the predicted number of target instances. This process can be formulated as:

### v4

# CoT-RVS: Zero-Shot Chain-of-Thought Reasoning Segmentation for Videos

@@ Abstract
Reasoning Video Object Segmentation is a challenging task, aiming at generating a mask sequence from an input video given a complex and implicit text query. While existing works finetune Multimodal Large Language Models (MLLM) for the task, they still fail in video inputs given complex temporally-sensitive queries, indicating their lack of temporal and spatial integration in complex scenarios. In this paper, we propose CoT-RVS, a novel framework employing the zero-shot Chain-of-Thought (CoT) capability of MLLM to address these complex challenges by temporal-semantic reasoning: CoT-RVS analyzes the visible objects within a given frame that possibly match the language query (semantic), and chooses a corresponding keyframe for each object that can be observed effortlessly among all frames (temporal). Notably, the CoT-RVS framework is training-free and compatible with closed-source MLLMs, which can be applied to Reasoning Video Instance Segmentation. Our framework’s training-free feature further allows its extension to process online video streams, where the CoT is used at test time to update the object of interest when a better target starts to emerge and becomes visible. We conduct extensive experiments on video object segmentation with explicit and implicit queries. The results show that CoT-RVS significantly outperforms previous works in both cases, qualitatively and quantitatively.
Figure 1: CoT-RVS is a novel framework for Reasoning Video Segmentation utilizing the zero-shot Chain-of-Thought (CoT) capability of pretrained multimodal Large Language Models, and correctly segments the fast-moving player right before, during, and after his three-point shot after CoT, given a temporally sensitive query “Which player makes a three-point shot in this basketball game?” which requires both spatial and temporal reasoning.
@@ 1 Introduction
Reasoning Video Object Segmentation (Reasoning VOS) presents a formidable challenge in the field of computer vision, requiring the generation of a mask sequence from an input video alongside an implicit and often complex text query (Yan et al., 2024; Bai et al., 2024). Unlike traditional vision-language tasks that directly associate visual data with textual descriptions, Reasoning VOS requires more advanced cognitive capabilities due to the dynamic nature of video data, where factors such as temporally sensitive queries, occlusions and disocclusions due to rapidly moving object can complicate the segmentation process.
@@
Existing approaches of reasoning segmentation focus on finetuning the multimodal Large Language Models (MLLMs) for segmentation-based tokens. While achieving good results in Reasoning VOS, the finetuning process is time consuming, and they still suffer from low accuracy when temporal reasoning is required. An implicit and time sensitive query, “Which player makes a three-point shot in this basketball game?”, for instance, is more challenging than “Which player is wearing black?” (Fig.˜1). A query is time or temporally sensitive if the relevant video objects with the specified action context (e.g., three-point shot), do not appear uniformly over time but rather are concentrated at certain intervals, requiring high-level temporal reasoning to correctly find them to the satisfaction of the query requirement. Even though reasoning segmentation has achieved impressive performance on images (Lai et al., 2023; Kao et al., 2025; Liu et al., 2025) and videos (Bai et al., 2024), current results are far from perfect in the video domain, showcasing failure in integrating temporal information with spatial and textual context in satisfying complex and time-sensitive queries.
To address this challenge, this paper aims at injecting stronger temporal reasoning ability into a segmentation framework through diversely pre-trained MLLMs empowered by zero-shot Chain-of-Thoughts (Wei et al., 2022). We enable new model capabilities through carefully designed task-specific prompts and innovative module interaction, a visionary trend also shared by Guo et al. (2022); Lian et al. (2023); Hu et al. (2023). Thus, we propose CoT-RVS, a novel, modular framework, whose key success lies in its exceptional reasoning capability on the temporal-semantic correlation across keyframe candidates, see an example in Fig. 2. At inference time, CoT-RVS analyzes the visible objects within a given frame that possibly matches the language query (semantic), and chooses a corresponding keyframe for each instance of interest, where the object can be observed with the least effort among all frames (temporal).
Thus, our CoT analysis requires complex, human-level reasoning within and among different video frames, e.g., “Frame 3 and 4 contains the player in shooting posture, which satisfies the user query, while frame 3 is a better keyframe with a clearer view.”. We will show in this paper that an accurate, in-depth reasoning for this temporal-semantic correlation will significantly improve the model performance in processing time sensitive query.
@@
One of the other standout features of the CoT-RVS framework is its training-free nature, making it compatible with open-source and closed-source MLLMs. We leverage the state-of-the-art Gemma3 (Team et al., 2025) and GPT-4o (Hurst et al., 2024) as the “thinking” agent in our CoT-RVS framework to generate keyframe selectivity. To evaluate the effectiveness of CoT-RVS, we conduct comprehensive experiments on referring VOS datasets (e.g., MeViS (Ding et al., 2023a) and Refer-DAVIS-17 (Pont-Tuset et al., 2017)) and reasoning VOS datasets (e.g., ReVOS (Yan et al., 2024) and ReasonVOS (Bai et al., 2024)). We demonstrate that CoT-RVS significantly outperforms existing referring and reasoning VOS approaches, qualitatively and quantitatively, with its more general formulation based on Reasoning Video Instance Segmentation (Reasoning VIS), which can output multiple instance-level mask sequences to satisfy the user query. In addition, we propose a variant version of CoT-RVS for online video streams, where CoT is used to update the target object when a new object starts to appear in the middle of the video that aligns better with the text query.
In summary, our contributions include:
  • A modular, zero-shot reasoning VOS framework that coordinates MLLMs, image segmenters, and video processors in both offline and online settings, without any training or fine-tuning, unlike prior methods.
  • A keyframe selection pipeline based on CoT prompting, specifically tailored for temporal reasoning, where MLLMs are required to both localize and describe scene-relevant frames, beyond simple object retrieval.
  • An online reasoning extension that adaptively re-selects keyframes at inference time, a feature rarely explored in existing R-VOS systems, enabling streaming video processing.
Figure 2: Keyframe selection is particularly challenging when processing time sensitive queries. Given a video, where two cats are playing with the cat teaser at the beginning, and then the white cat stays static in the second half of the video, existing methods e.g., VISA (Yan et al., 2024) fail to find a proper keyframe and directly use the user prompt for reasoning segmentation. In contrast, CoT-RVS extracts temporal-semantic correlation from the input video and successfully outputs a more reasonable keyframe with detailed description to the target object.
@@ 2 Related Work
Video segmentation. Video object segmentation (VOS) tracks objects in a given video and output corresponding mask sequences. Early VOS approaches (Maninis et al., 2018; Caelles et al., 2017) rely on online first-frame finetuning, thereby constrained to specific applications because of the high time cost. To speed up inference, follow-up works adopt efficient online learning algorithm (Robinson et al., 2020; Bhat et al., 2020), MRF graph inference (Bao et al., 2018), temporal CNN (Xu et al., 2019; Hou et al., 2019), capsule routing (Duarte et al., 2019), tracking (Jang and Kim, 2017; Wang et al., 2019; Perazzi et al., 2017; Chen et al., 2020), embedding learning (Yang et al., 2020; Voigtlaender et al., 2019; Chen et al., 2018) and space-time matching (Hu et al., 2018; Oh et al., 2019a; Cheng et al., 2021b). Recent VOS methods allow various control modalities, such as categories (Cheng et al., 2021a; Wang et al., 2021; Wu et al., 2022d; Zhang et al., 2023c; 2025; Zhou et al., 2024), points (Ravi et al., 2024; Rajič et al., 2025; Homayounfar et al., 2021), segmentation masks (Cheng and Schwing, 2022; Cheng et al., 2018; Oh et al., 2019b; Perazzi et al., 2017; Ravi et al., 2024), or text query (Botach et al., 2022a; Ding et al., 2023a; Wu et al., 2023b; Seo et al., 2020; Ye et al., 2019; Li et al., 2023b; Wu et al., 2022b; Miao et al., 2023a; Cuttano et al., 2024), making the segmentation more flexible and interactive.
@@
Specifically, building on the advanced development of LLM, Reasoning VOS focuses on generating mask sequences from a highly implicit and complex text query. VISA (Yan et al., 2024) first selects a keyframe and follows the core idea of LISA (Lai et al., 2023) by generating a segmentation-oriented token with finetuned image-based MLLM. Video-LISA (Bai et al., 2024) and HyperSeg (Wei et al., 2024) encode the video input in a multi-grained manner and concatenate the text query, feeding both to the LLM for segmentation-based outputs. However, none of the prior works incorporate adequate consideration of CoT reasoning ability for spatial and temporal information in Reasoning VOS. Our CoT-RVS is the first zero-shot-CoT framework compatible with pre-trained MLLM and image segmentation module, making it flexible to a boarder audience. In fact, recent studies, e.g. (Snell et al., 2024), have emphasized that test-time compute and inference-time reasoning are becoming increasingly central to leveraging pre-trained LLMs. Our method is aligned with this direction.
@@
Chain-of-Thought MLLM. Drawing inspiration from the exceptional reasoning capabilities of Large Language Models (LLMs), Multi-modal Large Language Models (MLLMs) seek to integrate visual components to enhance multimodal perception and reasoning. These models are often employed to produce captions or respond to inquiries based on visual and textual inputs, typically achieved by techniques such as prompt engineering (Liu et al., 2023c; Shen et al., 2023; Wu et al., 2023a), cross-attention mechanism (Alayrac et al., 2022), Q-Former (Dai et al., 2023; Li et al., 2023a), prompt tuning (Zhang et al., 2023b), linear projection (Koh et al., 2023; Liu et al., 2023b), and unified model architectures (Peng et al., 2023; Xiao et al., 2024). These MLLMs have greatly revolutionized the traditional vision-centric tasks like image segmentation (Lai et al., 2023; Xia et al., 2024; Zhang et al., 2023a; He et al., 2024; Bao et al., 2024), object detection (Zhang et al., 2023a; Zang et al., 2024; You et al., 2023; Chen et al., 2023; Ma et al., 2024), video segmentation (Yan et al., 2024; Bai et al., 2024; Wei et al., 2024), and object tracking (Wang et al., 2024; Munasinghe et al., 2023; Zhu et al., 2023).
@@
In addition to these advancements, CoT-based LLMs have been proposed to improve the reasoning capability (Wei et al., 2022; Wang et al., 2022). Existing CoT approaches tend to concentrate on inference through progressive reasoning demonstrations (Zhang et al., 2023d; Wei et al., 2022; Lyu et al., 2023) or by utilizing explicit prompts that clearly outline each step (Kojima et al., 2022). However, enabling CoT capabilities for MLLMs typically undergo fine-tuning on specifically designed multimodal CoT datasets (Mondal et al., 2024; Zhang et al., 2023e; Lu et al., 2022) or incorporate intricate intermediate representations (Mitra et al., 2024; Surís et al., 2023), which restricts their usability and scalability due to data accessibility and time complexity. Closer to our work, recent works, such as ThinkFirst (Kao et al., 2025) and Seg-Zero (Liu et al., 2025), have successfully incorporate CoT MLLM in reasoning image segmentation, while still falling short of “thinking” in the temporal domain. In contrast, we aim to, for the first time, abridge this gap by exploring zero-shot CoT reasoning of pre-trained MLLMs and apply them in the area of video segmentation.
3 CoT-RVS: Multi-Agent Framework
Figure 3: CoT-RVS for Reasoning VIS where Reasoning VOS is a special case.
@@ 3 CoT-RVS: Multi-Agent Framework
CoT-RVS is a multi-agent framework that consists of three network components: an MLLM keyframe selector $\mathcal{F}_{key}$, a reasoning image segmentation model $\mathcal{F}_{seg}$, and a video processor $\mathcal{F}_{vid}$, see Fig.˜3.
@@
Input and Output Settings. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$ that can be very complex, CoT-RVS is inherently a Reasoning Video Instance Segmentation model $\psi_{\theta}(\cdot)$ which generates a list of non-overlapping binary mask sequences $\hat{y}=\left\{\mathcal{M}_{i}\in\mathbb{R}^{T\times H\times W}\right\}_{i=1}^{k}$:
@@
where $k$ indicates the number of object instances of interest that satisfy the requirements of input query $q$, where Reasoning Video Object Segmentation (VOS) is clearly reducible to Reasoning Video Instance Segmentation (VIS) by setting $k=1$.
@@
Query requirement. CoT-RVS shares a similar query setting with the first Reasoning Video Object Segmentation (Yan et al., 2024), where the input query $q$ is complex or implicit (e.g., “Please find the vehicle in the video which can accommodate at least four people.”), in contrast to the Referring VOS (Seo et al., 2020) task which takes explicit and direct query (e.g. “Please segment the red car in the video.”).
@@ 3.1 Multi-Agentic Framework
Figure 4: Illustration of the CoT Process. The MLLM agent is prompted to synthesize chain-of-thought questions and answers. This CoT process enhances the in-depth temporal and spatial understanding for the keyframe candidates. Based on the CoT result, the MLLM agent outputs an instance-level object list, which contains the object-specific keyframe and description within the frame. Refer to the Appendices.˜A and E for the detailed prompt and output examples.
MLLM Keyframe selector. The keyframe selector is an MLLM agent, e.g., GPT-4o (Hurst et al., 2024) or Gemma3 (Team et al., 2025), that generates keyframe selectivites from a list of keyframe candidates, denoted as $\{I^{c}_{i}\}$, with a CoT thinking process that analyzes the temporal-semantic correlation among the input video frames. To find all the object instances that meet the query requirement necessitating temporal-semantic understanding, we select their corresponding keyframes such that the pertinent instance is most apparent for the MLLM agent to observe from that frame. Thus, the output contains a list of selected object instances, each of which corresponds to an object index, a respective keyframe, and a keyframe-related object description (e.g., “the player in black jersey at the right side of the frame”).
Refer to Appx.˜A for the prompting details. In implementation, the agent is prompted to automatically synthesize a series of questions for each keyframe candidate and answer them one-by-one. Starting from general semantic information like “What can be seen in the frame?”, to temporal-related questions like “Is it a better keyframe?”, and eventually ending with detailed questions like “Are there any new objects that meet the query requirement?”, this CoT process enables thorough reasoning between the user query and the temporal-semantic correlation among the keyframe candidates.
After analyzing all the keyframe candidates, the agent will collect all the per-frame responses and generate an output list composed of all target instances along with the respective keyframes based on its progressive question-and-answer pairs. We demonstrate this CoT process in Fig.˜4 for visualization.
@@
3.2 CoT-RVS: Reasoning Video Instance Segmentation
CoT can be applied to the entire video when available for reasoning instance segmentation. Given an input video $x_{v}=\left\{I_{i}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{T}$ and a text query $q$, we first uniformly sample a number of keyframe candidates $\{I^{c}_{i}\in x_{v}\}_{i=1}^{T^{\prime}}$, where $T^{\prime}=\lfloor T/\xi\rfloor$ for some $\xi<T$. These candidates, along with the user query $q$, are then sent to the MLLM keyframe selector $\mathcal{F}_{key}$ for temporal-semantic analysis, as described in Sec.˜3.1, which eventually outputs a keyframe selectivity $S$ composed of a list of instance-level text descriptions $\{s_{i}\}_{i=1}^{k}$ with respect to their corresponding keyframes $\{f_{i}\}_{i=1}^{k}$ (e.g., “the red car” in $I^{c}_{2}$, “the blue motorcycle” in $I^{c}_{4}$, …). Specifically, $k$ is the predicted number of target instances. This CoT process can be simplified as: